# Visuospatial navigation without distance, prediction, integration, or maps

### v3

@@ 3 Discussion
Cognitive maps, widely accepted as the ultimate way many organisms represent space, describe a system of cells tuned to distinct spatial elements collectively composing a Euclidean graph, upon which novel shortcuts can be calculated via simple vector transformation [27]. Response-based navigation, on the other hand, uses primarily visual observations to make movement decisions, enabling bottom-up construction of a more flexible non-Euclidean route topology or scene system [28, 29, 8, 9, 30]. However our understanding of egocentric, response-based strategies lags behind that of cognitive maps, in part due to historical preference for the latter [31, 32, 33, 34, 35]. Predispositions toward maps persist via experimental paradigms that reinforce the primacy of Euclidean graphs, even despite results demonstrating similar rates of navigational success with either strategy [6, 7, 36]. While flexible shortcut use or direct trajectories are certainly optimal in certain scenarios, overemphasis on cognitive maps denies both understanding and value of response-based systems.
@@
Given that map-based knowledge demands construction of a recurrent internal world model to predict and integrate external dynamics, complexity required presumably far surpasses systems that directly respond to visual input. Assuming a given organism would seek to minimize computational complexity while maintaining successful performance, external factors such as environmental context play a critical role in determining which system preferentially develops. For example, path integration may be ideal for desert ants which require precision to quickly return home after foraging in midday heat [37], though may cause problems in unpredictable atmospheres with potentially dangerous shortcuts. Response-based representations, in contrast, have been suggested to favor dense environments with many navigational constraints or without ambiguous landmarks such as jungles or mazes [38, 9, 22]. What has not been evidently studied is their sufficiency in open arenas, an environment suggested to favor cognitive maps [4]. Our results present non-intuitive evidence: navigational strategies based on visual responses can perform well in open arenas.
Map-based navigation often assumes a framwork based on distance prediction [3, 4, 5] or vestibular integration [1, 2] in order to construct a robust model. However, key visual dimensions were found to be functionally irrelevant for the two map-based algorithms that performed ablation studies [1, 3], suggesting that while vision may be translated into either framework, it provides a separable informational stream. Further signaling such degeneracy in information and resulting strategy, recent studies have shown that by restricting sensory or cognitive capability, rodents lose map-based spatial selectivity while maintaining flexible route navigability based on visual angles and sequential memory [39, 40]. A bottom-up approach aiming at identifying minimally sufficient processes for effective behavior remains valuable for providing the complementary perspective to studying range in potential capability.
@@
Rodent paths observed in the Morris Water Maze reflect movement characteristics from each of the three classes in the present study. While many learn how to take optimal direct paths, others travel in clearly wrong directions before abruptly turning toward the patch [41, 42, 43]. Although often termed "indirect search", this pattern parallels an indirect sequential navigational strategy. Other rodents instead walk toward the patch yet spin as they approach, attributed as "self-orienting" or even "nonsense movements" of "directed search" [44, 41, 42, 43], they mirror biased diffusive trajectories. Shifting perspective to view such behavior as essential components to response-based routes, as opposed to search behavior or errors relative to ideal trajectories, may yield novel hypotheses. For example, the balanced relative occurence metrics and across-trial stability of the three strategies [44, 41] suggest each may confer contextual tradeoffs, where diversity has evolved to ensure robust population-level fitness in a variable environment.
@@
The biased diffusive trajectories appear to mirror the helices of sperm cell chemotaxis [45, 46]. Sperm cells modulate swimming path curvature via temporally sampled chemical concentrations to bias motion up-gradient towards an egg. Despite clear differences in perception and movement capabilities, both learn a similar movement strategy to robustly navigate to a hidden target, suggesting fundamental properties for gradient navigation, which may be either directly sampled or indirectly constructed via angle-based vision.
@@
Elliptical decision-making thresholds have been recently found to govern spatial bifurcations between choices or goal vectors [47, 48], in the same way ellipses spanning adjacent corners govern turn decisions for the indirect sequential agents. Fruit flies, desert locusts, and zebrafish have each been observed to move toward the average of two choices until reaching a critical threshold, spontaneously forcing a decision towards one or the other choice, even in sequence if more than two are available. While in the current study, perception-action loops learned to deterministically use elliptical thresholds to compose route segments, here they were used by a ring attractor model to stochastically choose between goals or sets of goals. Despite differing computational, algorithmic, and implementational level details, resulting behavior appears to have converged on a fundamental element: angle-based turn decisions with respect to two spatially distinct entities.
@@
Rather than assuming the entire range of perceptual and cognitive capabilities of organisms when modeling behavior, probing minimal sufficient conditions for successful behavior under energetic-bound constraints offers unique insight into how organisms may behave outside the lab. Though many organisms may have the ability to perceive depth, construct predictive frameworks, and integrate vestibular cues, a simpler feedforward system might do just fine, without the energy expenditure. That the resulting behavior tracks with at least three separate phenomena covering rodents, sperm cells, insects, and fish provides evidence that such constraints may be widely applicable.
This is not intended to argue against the importance for cognitive maps, prediction, distance perception, or vestibular integration, nor to deny crossmodal interactions between map and reponse pathways, but rather encourage a plurality of perspectives moving forward. Given our historical bias toward top-down representation, as well as our physical experience using maps or viewing rodents in a maze beneath us, we need to approach the egocentric, response-based perspective with a new lens. Response-based navigation can provide a fully separable navigational strategy with its own mechanisms and logic, demanding novel experimental methods for disentangling systems and characterizing mechanisms, proposing alternative navigational hypotheses to contend with assumptions of search or errors. Each system has characteristic advantages and disadvantages, with evolution and development shifting the balance toward a particular direction in a given environmental context. Only by understanding both can we more fully grasp the breadth of tools potentially available to a navigating organism.
4 Methods
@@
Vision is simulated as a raycast originating from a single retina at the front edge of the agent, the orientation of both visual retina and agent are bound as a single state variable. Perceptual input is structured as a 1D array with each element one-hot encoded with respect to wall identity. Though a simplification from 2D vision, collapsing the vertical dimension has been previously found to retain sufficient directional information for visual navigation [49, 50].
Two key hyperparameters govern the raycasting process: field of vision (FOV) and resolution. FOV stretches the limits of the raycast with respect to the agent, and resolution determines the number of rays to cast within these limits. The FOV is set at 40% or 144 degrees to mirror the functional visual field for humans, as measured by reaction time to discrimination tasks [51], though the impact of FOV on navigation performance was also assessed (Fig. S1 1st row). The visual resolution ($\upsilon$) is minimally initialized at 8 rays for much of the analysis and later varied to examine its effect on behavior (Fig. S1 2nd row).
Weber-Fechner logarithmic scaling is applied for distance information, similar to how height is perceived by our eyes [52]. The scaling is calculated as follows:
@@
The arena is defined at 1000 x 1000 units, the patch at 50 units, the maximum speed of the agent at 2 units per timestep. This may correspond either to 15 cm/s swimming speed, a 3.75 m patch radius, and a 75 x 75 m arena for rodents, or 1.5 m/s walking speed, a 37.5 m patch radius, and a 750 x 750 m arena for humans. These sizes are significantly scaled up from the 10-15 cm diameter platform and 1-2 m diameter arena often used for rodents in the Morris Water Maze [44, 41, 42, 43, 1]. Assuming an average laboratory environment with a width of about 5-15 m, the simulated boundary corners may be classified as exceptionally distal landmarks, and the environment as a very open arena. Our focus on distal landmarks runs contrary to the algorithm trained by Banino et al, where ablation studies demonstrate such landmarks to be functionally irrelevant [1].
Environmental geometry provides the spatial reference needed for both orientation and location [53]. Although landmark and geometric encoding have been argued to be performed by separate modules [54, 55, 56, 57], in the present study the two are not differentiated.
@@
The CNN used is a simplified version of the ConvNeXt v2 architecture [58, 59], a state-of-the-art design that outperforms earlier varitions in the CNN design space and rivals the best vision transformers. While simpler designs exist, the ConvNeXt was chosen for its separable depthwise and pointwise convolutions, where separate parameters act on the orthogonal depth and channel axes, enabling greater expressibility for relatively low additional computational cost. A transformer was not chosen since they lack biological inductive biases, such as translation equivariance, and due to their novelty, lack the understanding and tools currently available for CNNs [60].
Beyond critical minimal values, depth and dimension of the CNN and perceptron did not noticeably affect performance (Fig. S1 3rd/4th rows). Notably, two channel outputs for the CNN is insufficient for agents to reliably solve the task, possibly reflecting an intrinsic, minimal dimensionality to the navigation problem [61]. For more complex and recurrent architectures, as well as those with proprioceptive feedback, data at adequate sample size is forthcoming at the time of submittal and will be appended upon review. The activation function used for the CNN is a rectified linear unit. A sigmoid linear unit was found to net marginal performance increases, though slower simulation times.
The number of parameters is less than 300 for the models presented in the main text, significantly less than the tens of thousands needed for the predictive navigation network used in [3], and still less than the thousand needed for the non-visual linearized predictive networks in [5].
@@
Network output is a single continuous value that directly represents turning angle for the next timestep, bounded by 90°90°90  ° left to 90°90°90  ° right via a hyperbolic tangent function. Speed is scaled by the output via a linear function, with zero representing maximum speed and both left and right limits for when the agent is stopped.
@@
Evolutionary Algorithm
The ES algorithm used is Policy Gradient Parameter Exploration (PGPE) using the ClipUp optimizer [62, 63], with parameters outlined in Table 1. Among various ES algorithms, PGPE was chosen for its balance between performance and speed, with greater performance than OpenAI-ES and greater speed than CMA-ES [64]. ES entails less complexity than RL, as there is no need for differentiability, value approximation, or within-trial credit assignment [25, 26]. The main drawback of ES, sample inefficiency, is acceptable given its parallelizability.
@@
Fitness performance during training was calculated as the time taken to reach the patch plus remaining distance. Alternative scalings were tried but found to not make a significant difference.
@@
The agent was tested by simulating the top agent of an evolutionary run at different intializations across the grid, spaced every 25 units and oriented in 16 directions. Each initialization was run for 500 timesteps, regardless if the agent collides with the patch. The test data from this set of trajectories was used for Fig. 2.

### v4

@@ 3 Discussion
We trained a simple yet expressive artificial neural network to navigate toward a hidden patch in a robust and efficient manner (Fig 1, S1). By placing heavy constraints on perception, cognition, movement, and the environment, that is, monocular 1D visual perception, no predictive framework or within-trial memory, a single decision variable governing movement kinematics (turning & speed), and a four wall or landmark environment, we demonstrate the sufficiency of an egocentric, feedforward, route-based framework to solve the Morris Water Maze.
Three distinct classes of algorithms emerged in order to accomplish this, each with a set of distinguishable features (Fig 2, 3). The first class of agents navigate by composing a sequence of straight segments, indirectly oriented with respect to the patch, with major turns defined by elliptical-shaped manifolds, i.e. view-specific invariances stemming from perception of two adjacent corners. The second class instead learns to spin or diffuse toward the patch, driven by smoother global gradients constructed from the same elliptical invariances. And the third uses distance perception, a phenomenological implementation of advanced perceptual ability, to calculate near-ideal paths to the patch. Described via behavioral characteristics, these classes are termed indirect sequential, biased diffusive, and direct pathing, respectively.
The classes can be separated by two correlation metrics: decorrelation time and directedness (Fig 2), indicating temporal persistence in heading and directional potential in space. Biased diffusive has low measures for both, indirect sequential has medium, and direct pathing has high. The three classes tend to cluster into distinct bins in this 2D phase space, suggesting fitness-based attractors for movement correlation parameters.
Angle-based algorithms (indirect sequential and biased diffusive) confer comparable performance though prefer different visual resolutions (Fig 4). Indirect sequential agents evolve more readily and perform better with lower visual resolution and biased diffusive with higher. While only requiring low quality visual signal and sparse orientational profiles, the indirect sequential agents are highly susceptible to visual angle noise throwing them off-course, whereas such noise does not affect the dynamic, disperse process of biased diffusive agents. Distance-based, direct pathing agents perform better than either angle-based class, resulting in an abrupt phase transition as the distance input becomes computationally usable, although requiring another dimension to their perception.
@@
Our understanding of route-based strategies lags behind that of cognitive maps, in part due to historical preference for the latter hartDevelopmentSpatialCognition1973; siegelDevelopmentSpatialRepresentations1975; warrenNonEuclideanNavigation2019; wangEgocentricAllocentricRepresentations2020; kunzNeuralCodeEgocentric2021. Predispositions toward maps persist via experimental paradigms that reinforce the primacy of Euclidean graphs, despite partial dissociability and parallel use of the two strategies, and even despite results demonstrating similar rates of navigational success with either strategy marchetteCognitiveMappersCreatures2011; sheltonChapterSixMechanistic2013; weisbergCognitiveMapsPeople2018; goodroeComplexNatureHippocampalStriatal2018. Whereas calculating shortcuts or direct trajectories are certainly optimal in certain scenarios, overemphasis on cognitive maps denies both value and understanding of route-based systems.
@@
Given that map-based knowledge demands construction of a recurrent internal world model to predict and integrate external dynamics, complexity required presumably far surpasses systems that directly respond to visual input. Assuming a given organism would seek to minimize computational complexity while maintaining successful performance, external factors such as environmental context play a critical role in determining which system preferentially develops. For example, path integration may be ideal for desert ants which require precision to quickly return home after foraging in midday heat wehnerParallelEvolutionThermophilia2011, though may cause problems in unpredictable areas with potentially dangerous shortcuts. Route-based representations, in contrast, have been suggested to favor dense environments with many navigational constraints or without ambiguous landmarks such as jungles or mazes babayanHippocampocerebellarCentredNetwork2017; peerStructuringKnowledgeCognitive2021; spiersHowEnvironmentShapes2022. What has not been evidently studied is their sufficiency in open arenas, an environment suggested to favor cognitive maps kesslerHumanNavigationStrategies2024. Our results present non-intuitive evidence: navigational strategies based on visual routes can perform well in open arenas.
The reinforcement learning community represents the navigational dipole somewhat differently. Model-based learning entails planning upon a predictive, goal-directed framework by constructing a model of external dynamics, while model-free works directly on cached action-value pairs dawUncertaintybasedCompetitionPrefrontal2005; woodPsychologyHabit2016. Described in this way, model-based learning appears to reflect map-based strategies and model-free to response sequences, coordinated through associative memory, babayanHippocampocerebellarCentredNetwork2017; anggrainiNeuralSignaturesReinforcement2018, though the two perspectives may not fully align khamassiIntegratingCorticolimbicbasalGanglia2012; herwegSpatialRepresentationsHuman2018; vijayabaskaranNavigationTaskAction2022. Despite discrepancies, balanced theoretical and experimental effort toward both model-based and model-free reveals contextual advantages for each. Model-based learning tends to be more efficient in learning, as well as flexible and theoretically effective in control, yet require significantly higher capacity to handle complex external dynamics, which in turn may lead to overfitting nagabandiNeuralNetworkDynamics2018; kimTaskComplexityInteracts2019. In contrast, model-free learning is computationally cheap and effective at scale.
Neural connections
The hippocampus and striatum are typically contrasted with respect to the Morris Water Maze, where the former is responsible for map-based, allocentric strategy with a focus on environment boundaries and distal (extra-maze) landmarks, and the latter for response-based, egocentric strategy using proximal (intra-maze) landmarks doellerParallelStriatalHippocampal2008; goodroeComplexNatureHippocampalStriatal2018; parra-barreroMapSpatialNavigation2023. One possibility not examined in this dual-solution focus is the construction of egocentric routes based on boundaries and distal landmarks harveyEmergenceEgocentricCue2008. Accounting for this third option, i.e. equating routes as sequences of decisions, may help unite seemingly disparate functions of the hippocampus: spatial navigation and episodic memory goodroeComplexNatureHippocampalStriatal2018; wangEgocentricCodingExternal2018; wangEgocentricAllocentricRepresentations2020; kunzNeuralCodeEgocentric2021; mooreLinkingHippocampalMultiplexed2021; parra-barreroMapSpatialNavigation2023. Rodents with hippocampal lesions are confined to beaconing, thigmotaxis, and scanning behavior riceLesionsHippocampusDorsolateral2015; higakiRecognitionEarlyStage2018; curdtSearchStrategyAnalysis2022; parra-barreroMapSpatialNavigation2023, while this can be understood as inflexible striatal navigation, it does not relate to the robust, initialization-invariant routes seen in our trained agents. The neural-semantic dipole between hippocampal-maps and striatal-responses does not leave room to allow for flexible route-based navigation, which presumably uses hippocampal episodic sequences to navigate via both egocentric responses and allocentric boundary or distal information, yet without demanding the ability to localize or plan on a map-like substrate.
The distributed nature of these systems generates degeneracies on multiple fronts in the solution space, allowing an organism to choose which pathway might suit the context. Outside the hippocampal-striatal dipole, several other cortical areas process navigational information, including: the retrosplenial cortex augerRetrosplenialCortexCodes2012, occiptal parietal cortex bonnerCodingNavigationalAffordances2017, and even visual cortex saleemCoherentEncodingSubjective2018. As described in the introduction, two recent models that allow for visual information, as well as vestibular baninoVectorbasedNavigationUsing2018 or distance recanatesiPredictiveLearningNetwork2021, did not learn visual dependencies. While the former two information streams and learned representations are sufficient for the task, they are incomplete on their own parra-barreroMapSpatialNavigation2023; krausseHeaddirectionSystemShows2025. Reflected in neural architecture, two distinct sets of head direction cells are suggested to be updated via either self-motion or visual cues jacobIndependentLandmarkdominatedHeaddirection2017; kornienkoNonrhythmicHeaddirectionCells2018, though see also evidence of integration siegenthalerVisualObjectsRefine2025. Guidance and aiming tasks have been shown to prefer allocentric and egocentric action spaces, respectively, though both can be solved with the other regardless vijayabaskaranNavigationTaskAction2022. An important point not often discussed, let alone remedied, the Morris Water Maze does not adequately separate guidance from map navigation parra-barreroMapSpatialNavigation2023: since landmarks are available from every point in the environment, triangulation via visual angles may be easier than self-localizing and computing vectors on a coordinate system wolbersChallengesIdentifyingNeural2014; eichenbaumRoleHippocampusNavigation2017. Degeneracy in the cognitive, sensory, action, and task spaces affords no easy answers. One way to inspect such systems is to focus on silencing or perturbing parts of these pathways riceLesionsHippocampusDorsolateral2015; aghajanImpairedSpatialSelectivity2015; acharyaCausalInfluenceVisual2016; higakiRecognitionEarlyStage2018; mooreLinkingHippocampalMultiplexed2021; vijayabaskaranNavigationTaskAction2022, and another, used here, is to explore a simple model constrained along one particular, under-studied pathway.
But does the agent learn a map? The agent learns to approximate a function that transforms egocentric observations into useful actions. Learned neural representations show significant directional modulation (Fig 3), aligning with a flight of papers on rodents in VR aghajanImpairedSpatialSelectivity2015; acharyaCausalInfluenceVisual2016; mooreLinkingHippocampalMultiplexed2021 and a modeling study examining task and action constraints vijayabaskaranNavigationTaskAction2022. Neural tuning corresponds with patch location, mirroring reports of representations with respect to goals sarelVectorialRepresentationSpatial2017; ormondHippocampalPlaceCells2022, as well as boundaries and landmarks wangEgocentricCodingExternal2018; alexanderEgocentricBoundaryVector2020; siegenthalerVisualObjectsRefine2025. The elliptical manifolds of the indirect sequential strategy trace view-specific correlations widely observed in primates and humans rollsSpatialViewCells1997; ekstromCellularNetworksUnderlying2003; jacobsSenseDirectionHuman2010; maoSpatialModulationHippocampal2021; pizaPrimacyVisionShapes2024, which may be linked to egocentric cells that conjunctively encode distance kunzNeuralCodeEgocentric2021. While we do not rule out the possibility of adapting these learned visuospatial functions to a map where the agent can plan trajectories, through the architectural constraints of our model, we have demonstrated that this externalization is not necessary. Similar to a minimal model by Bastien & Romanczuk proving vision alone can explain collective behavior without the need for higher level representations bastienModelCollectiveBehavior2020, and echoed by studies on sensorimotor contingencies and enactivist robotics hoffmannRobotsPowerfulAllies2018; degenaarSensorimotorTheoryEnactivism2017; egbertUsingEnactiveRobotics2022, the main point of our model is to show that directly perceived routes provide sufficient navigation without requiring map-based inference.
Moreover, correlating average neural activity with the external environment may mislead. The averaging of neural activity of our evolved agents over space and time reveals activity patterns correlated with goal direction (Fig 3A/B, right). However, a corresponding goal signal is absent in time-resolved activation patterns governing actual movement decisions of our agents (Fig 3D/E, right). Thus, the general question emerges whether empirically observed, averaged activation patterns correlating with higher-order task features are indeed involved in decision processes in real-time, or whether they may emerge as a "by-product" of learning the task via low-level decision rules (e.g. indirect sequential) and a corresponding averaging procedure. Or to put it in other words: to what extent are map-based representations, extracted through averaged activity patterns, directly involved in the generation of behavior, rather than a necessary statistical consequence of a complex neuronal system learning a spatial task? A recent studies shows how neural activity describes internal action sequences rather than external variables zutshiHippocampalNeuronalActivity2024, which may help explain why it is difficult for us to integrate between routes weisbergVariationsCognitiveMaps2014; warrenNonEuclideanNavigation2019. Characterizing behavior provides a complementary perspective to mean field neural analysis harveyEmergenceEgocentricCue2008, which may even be critical to understand direct perception-action behavior. While directional modulation may describe average neural correlation, clockwise curving trajectories may indicate a biased diffusive navigational strategy mooreLinkingHippocampalMultiplexed2021 or attractor-like alignment may indicate the use of visual manifolds vijayabaskaranNavigationTaskAction2022.
@@
Rodent paths observed in the Morris Water Maze reflect movement characteristics from each of the three classes in the present study. While many learn how to take optimal direct paths, others travel in clearly wrong directions before abruptly turning toward the patch cookePathfinderOpenSource2020; curdtSearchStrategyAnalysis2022; villarreal-silvaAgedRatsLearn2022. Although often termed "indirect search", this pattern parallels the indirect sequential navigational strategy. Other rodents instead walk toward the patch yet spin as they approach, attributed as "self-orienting" or even "nonsense movements" of "directed search" grazianoAutomaticRecognitionExplorative2003; cookePathfinderOpenSource2020; curdtSearchStrategyAnalysis2022; villarreal-silvaAgedRatsLearn2022, they mirror biased diffusive trajectories. Shifting perspective to view such behavior as essential components to egocentric navigational routes, as opposed to search behavior or errors relative to ideal trajectories, may yield novel hypotheses. Perhaps these rodents prioritize angular over distance information etienneNavigationSmallMammal1992; wallaceRatsMaintainOverhead2013; hagbiHeightsPlainsHow2022. Balanced relative occurence metrics and across-trial stability of the three strategies grazianoAutomaticRecognitionExplorative2003; cookePathfinderOpenSource2020 suggest each may confer contextual tradeoffs, where diversity has evolved to ensure robust population-level fitness in a variable environment.
@@
The biased diffusive trajectories appear to mirror the helices of sperm cell chemotaxis alvarezComputationalSpermCell2014; rodeInformationTheoryChemotactic2024. Sperm cells modulate swimming path curvature via temporally sampled chemical concentrations to bias motion up-gradient towards an egg. Despite clear differences in perception and movement capabilities, both learn a similar movement strategy to robustly navigate to a hidden target. Though clearly the mechanisms differ, as sperms cannot see, the higher level phenomena seem to overlap, suggesting fundamental algorithmic properties for gradient navigation, which may be either directly sampled or indirectly constructed via angle-based vision.
@@
Elliptical decision-making thresholds have been recently found to govern spatial bifurcations between choices or goal vectors sridharGeometryDecisionmakingIndividuals2021; gorbonosGeometricalStructureBifurcations2024, in the same way ellipses spanning adjacent corners govern turn decisions for the indirect sequential agents. Fruit flies, desert locusts, and zebrafish have each been observed to move toward the average of two choices until reaching a critical threshold, spontaneously forcing a decision towards one or the other choice, even in sequence if more than two are available. While in the current study, perception-action loops learned to deterministically use elliptical thresholds to compose route segments, here they were used by a ring attractor model to stochastically choose between goals or sets of goals. Despite differing computational, algorithmic, and implementational level details, resulting behavior appears to have converged on a fundamental element: angle-based or view-specific turn decisions with respect to two spatially distinct entities, proposing a non-Euclidean spatial representation.
While there are speculative comparisons to make between our model and biology, there are limitations. The model is intended to be phenomenological, thus only general relations should be made. Although our convolutional neural network may reflect the visual cortex lindsayConvolutionalNeuralNetworks2021, its operations as well as the downstream layers are meant to ambiguously define sensorimotor perception-action information flow rather than specific brain regions or operations. Complexity has been reduced to a simple feedforward network, whereas the visual cortex and associated structures contain recurrent connections. Similarly, neural network optimization accumulates thousands of parallel perturbations and experiences each round, and thus is intended to reflect neither the speed nor mechanism of biological learning, nor does it reflect the search process preceding navigation. Only the behavior of a fully trained network is intended to be related to behavior of fully trained animals or humans.
@@
Rather than assuming the entire range of perceptual and cognitive capabilities of organisms when modeling behavior, probing minimal sufficient conditions for successful behavior under energetic-bound constraints offers unique insight into how organisms may behave outside the lab. Though many organisms may have the ability to construct predictive frameworks or integrate vestibular cues, a simpler feedforward system might do just fine, without the energy expenditure. That the resulting behavior and neural correlates track with phenomena covering rodents, insects, fish, and primates, provides evidence that such constraints may be widely applicable. These simple models do not assume navigation behavior to solely rely on one of the observed strategies, but rather suggest low dimensional visual techniques that can be used to save memory or complexity.
This is not intended to argue against the importance for cognitive maps, distance perception and prediction, or vestibular integration, nor to deny crossmodal interactions between map and route pathways, but rather encourage a plurality of perspectives moving forward. Given our historical bias toward top-down representation, as well as our physical experience using maps or viewing rodents in a maze beneath us, we need to approach the egocentric, route-based perspective with a new lens. Route-based navigation can provide a fully separable navigational strategy with its own mechanisms and logic, demanding novel experimental methods for disentangling systems and characterizing mechanisms, proposing alternative navigational hypotheses to contend with assumptions of search or errors. Each system has characteristic advantages and disadvantages, with evolution and development shifting the balance toward a particular direction in a given environmental context. Only by understanding both can we more fully grasp the breadth of tools potentially available to a navigating organism.
4 Materials and methods
@@
Vision is simulated as a raycast originating from a single retina at the front edge of the agent, the orientation of both visual retina and agent are bound as a single state variable. Perceptual input is structured as a 1D array with each element one-hot encoded with respect to wall identity. Though a simplification from 2D vision, collapsing the vertical dimension has been previously found to retain sufficient directional information for visual navigation wystrachHowFieldView2016; zeilVisualNavigationProperties2023.
Two key hyperparameters govern the raycasting process: field of vision (FOV) and resolution. FOV stretches the limits of the raycast with respect to the agent, and resolution determines the number of rays to cast within these limits. The FOV is set at 40% or 144 degrees to mirror the functional visual field for humans, as measured by reaction time to discrimination tasks sandersAspectsSelectiveProcess1970, though the impact of FOV on navigation performance was also assessed (S1 Fig, 1st row). The visual resolution ($\upsilon$) is minimally initialized at 8 rays for much of the analysis and later varied to examine its effect on behavior (S1 Fig, 2nd row).
Weber-Fechner logarithmic scaling is applied for distance information, similar to how height is perceived by our eyes buzsakiLogdynamicBrainHow2014. The scaling is calculated as follows:
@@
The arena is defined at 1000 x 1000 units and the patch at 50 units. The patch to arena ratio correlates a 1-2 m diameter arena and 10-15 cm diameter platform often used for rodents in the Morris Water Maze grazianoAutomaticRecognitionExplorative2003; cookePathfinderOpenSource2020; curdtSearchStrategyAnalysis2022; villarreal-silvaAgedRatsLearn2022; baninoVectorbasedNavigationUsing2018. An arbitrary timescale was chosen for our simulations, with a maximum agent speed at 2 units per timestep. This could be shifted to describe a lower or greater decision and action speed to reflect rodents, humans, or other organisms.
Environmental geometry provides the spatial reference needed for both orientation and location chanObjectsLandmarksFunction2012. Although landmark and geometric encoding have been argued to be performed by separate modules wangHumanSpatialRepresentation2002; leeTwoSystemsSpatial2010; doellerDistinctErrorcorrectingIncidental2008; doellerParallelStriatalHippocampal2008, in the present study the two are not differentiated.
@@
The CNN used is a simplified version of the ConvNeXt v2 architecture liuConvNet2020s2022; wooConvNeXtV2CoDesigning2023, a state-of-the-art design that outperforms earlier varitions in the CNN design space and rivals the best vision transformers. While simpler designs exist, the ConvNeXt was chosen for its separable depthwise and pointwise convolutions, where separate parameters act on the orthogonal depth and channel axes, enabling greater expressibility for relatively low additional computational cost. A transformer was not chosen since they lack biological inductive biases, such as translation equivariance, and due to their novelty, lack the understanding and tools currently available for CNNs khanTransformersVisionSurvey2022.
Beyond critical minimal values, depth and dimension of the CNN and perceptron did not noticeably affect performance (S1 Fig, 3rd/4th rows). Notably, two channel outputs for the CNN is insufficient for agents to reliably solve the task, possibly reflecting an intrinsic, minimal dimensionality to the navigation problem liMeasuringIntrinsicDimension2018. For more complex and recurrent architectures, as well as those with proprioceptive feedback, data at adequate sample size is forthcoming at the time of submittal and will be appended upon review. The activation function used for the CNN is a rectified linear unit. A sigmoid linear unit was found to net marginal performance increases, though slower simulation times.
The number of parameters is less than 300 for the models presented in the main text, significantly less than the tens of thousands needed for the predictive navigation network used in recanatesiPredictiveLearningNetwork2021, and still less than the thousand needed for the non-visual linearized predictive networks in stocklLocalPredictionlearningHighdimensional2024.
@@
Network output is a single continuous value that directly represents turning angle for the next timestep, bounded by 90° left to 90° right via a hyperbolic tangent function. Speed is scaled by the output via a linear function, with zero representing maximum speed and both left and right limits for when the agent is stopped.
@@
Optimization
The ES algorithm used is Policy Gradient Parameter Exploration (PGPE) using the ClipUp optimizer sehnkeParameterexploringPolicyGradients2010; tokluClipUpSimplePowerful2020, with parameters outlined in Table 1. Among various ES algorithms, PGPE was chosen for its balance between performance and speed, with greater performance than OpenAI-ES and greater speed than CMA-ES haVisualGuideEvolution2017. ES entails less complexity than RL, as there is no need for differentiability, value approximation, or within-trial credit assignment salimansEvolutionStrategiesScalable2017; majidDeepReinforcementLearning2021. The main drawback of ES, sample inefficiency, is acceptable given its parallelizability.
Table 1: Hyperparameters used for the ES algorithm
Name Value
Generations 1000
Episodes 20
Population Size 50
Standard Deviation, Initial 0.1
Standard Deviation, Learning Rate 0.1
Standard Deviation, Max Change 0.2
Mean, Learning Rate 0.2
ClipUp, Momentum 0.8
ClipUp, Max Speed 0.4
@@
Fitness performance during training was calculated as the time taken to reach the patch plus remaining distance if the agent has not reached the patch within simulation time limit (1000 timesteps). Alternative scalings were tried but found to not make a significant difference.
@@
The agent was tested by simulating the top agent of an evolutionary run at different intializations across the grid, spaced every 25 units and oriented in 16 directions. Each initialization was run for 500 timesteps, regardless if the agent collides with the patch. The test data from this set of trajectories was used for Fig 2.